### v2

# Towards a unitary formulation of quantum field theory in curved space-time: the case of Schwarzschild black hole

Abstract
We argue that the origin of unitarity violation and information loss paradox in our understanding of black holes (BH) lies in the standard way of doing quantum field theory in curved space-time (QFTCS), which is heavily biased on intuition borrowed from classical General Relativity. In this paper, with the quantum first approach, we formulate a so-called direct-sum QFT (DQFT) in BH space-time based on a novel formulation of discrete space-time transformations in gravity that potentially restores unitarity. By invoking the quantum effects associated with the gravitational backreaction, we show that the Hawking quanta emerging outside of the Schwarzschild radius ($r_{S}=2GM$) cannot be independent of the quanta that continue to be inside $r_{S}$. This enables the information to be carried by Hawking quanta, but in the BH DQFT formalism, we do not get any firewalls. Furthermore, DQFT leads to the BH evaporation involving only pure states. This means the quantum mechanical effects at the BH horizon produce two components of a maximally entangled pure state in superselection sector Hilbert spaces. This construction enables pure states to evolve into pure states, restoring unitarity and observer complementarity. Finally, we discuss how our framework leaves important clues for formulating a scattering matrix and probing the nature of quantum gravity.
1 Introduction
Ever since Hawking realized black holes (BHs) are unstable quantum mechanically, the world of theoretical physics has taken a murky turn with conceptual conundrums such as unitarity and information paradoxes [1, 2, 3, 4, 5, 6]. Even classically, according to general relativity (GR), BHs are the most fascinating objects in the Universe which awe-inspire us to investigate and speculate. The former is arduously difficult but the latter is a less complicated task because it is based on our natural human ’intuition’/’imagination’ emanating from our extreme enthusiasm about understanding BHs. Our indelible quest for quantum gravity in the last decades has been most often doomed by the interference of our natural imagination and speculations [7, 8, 9, 10, 11, 12]. The main aim of this paper is to go very meticulously through the popularly ’accepted’ calculation by Hawking [1] in quantizing a Klein-Gordon (KG) field in the ’near-horizon’ space-time of Schwarzschild black hole (SBH). We embrace the well-known fact that ’time’ is a parameter in quantum theory and one must separate it from the ’time’ as we know from GR. In BH physics, we point out that the origin of unitarity violation and information paradox lies in the way notions of time are treated in quantization. To be more precise, in the original Hawking treatment of quantum field theory in curved space-time (QFTCS), classical intuitions regarding the arrow of time according to GR are put forward before understanding quantum fields in curved space-time. But we note that the direction of exploring nature is from quantum to classical limit rather than the opposite. Thus, we need the quantum first approach (QFA) which puts forward the rules of quantum theory before our classical understanding of space-time [13]. It is worth remembering that the arrow of time in quantum physics emerges from the way we impose initial and final conditions on the quantum states [14]. Thus, classically when we see an arrow of time we must understand it as the thermodynamic arrow of time. Since we carefully go through the conceptual jurisdictions of quantum theory in understanding (quantum) BHs, we request our readers to avoid any preconceived notions fixed in the literature, textbooks, and also popular science articles.
As it was elaborately discussed in a previous paper [15], the standard quantum field theory QFT (SQFT) in Minkowski spacetime is very well routed in understanding discrete space-time symmetries $\mathcal{C}\mathcal{P}\mathcal{T}$. SQFT has led us to one of the most important symmetries of nature, which is $\mathcal{C}\mathcal{P}\mathcal{T}$. This means that to achieve a consistent unitary QFTCS; we need to take a step back to identify first the discrete space-time symmetries of the curved manifold, which tells us how we must move forward with our quantization. Furthermore, as it was discussed in our earlier work, a consistency check for a unitary QFTCS is that at sufficiently short distance scales, i.e., much below the horizon scales of curved space-time (still surely very far away from Planck scales), we must recover SQFT as we know from Minkowski space-time. This means any S-matrix we formulate in curved space-time should be reduced to the SQFT S-matrix in sufficiently short distance scales. This can be stated in another way if we consider very high energy modes whose wave number powers are much larger than the background curvature invariants (at the horizon scale) then we expect the effects of curved space-time to disappear and we expect the SQFT to hold. This means, in colloquial wording, high energy modes do not see the curvature of space-time, and they behave like in Minkowski space-time. Applying this logic in the context of SBH brings a question as the curvature invariants diverge near the singularity at $r=0$. We stress that one really needs to work with spherical harmonics (with the Kruskal-Szekeres metric) to understand quantum fields in BH backgrounds. Hawking original paper was about quantizing the KG field in Schwarzschild space-time by making heavy use of spherical harmonics under the assumption that the background nature of space-time remains almost the same under perturbations. Moreover, the effect of curved space-time occurs for the modes whose wavelengths are comparable to that of the Schwarzschild horizon radius. If, at these basic levels, we see a unitarity violation problem, which is the fact that there will be an entangled mixed state that can be lost beyond the horizon, resulting in the observer’s universe with mixed states. This is called the situation of pure states evolving into mixed states (unitarity violation). It is a legitimate quest to save unitarity in the low-energy effective theory, which is just GR and quantum mechanics (together, they should lead to a consistent QFTCS). Thus, to clarify our agenda in the rest of the paper, we restrict ourselves to refining some steps of Hawking’s original paper of 1975 [1] and forget for the moment about all developments and the discussion of various possibilities proposed in the name of quantum gravity approaches. We request the respected reader to stay with us in our hunt for the resolution of paradoxes in BH (quantum) physics at the fundamental level within perturbative QFTCS. Our assumption is that the solution to the unitarity violation must be found in the initial approach of Hawking calculation [^1] Of course, needless to say, we totally acknowledge Hawking’s amazing efforts of opening a door into the world of (quantum) BHs back in that time. Hawking has shown nature’s most perplexing objects to be unstable quantum mechanically. Still, the paradoxes that emerged later on. Therefore, it is essential to ask critical questions. We thus offer a new understanding of the nature of Hawking radiation beyond those notions widely discussed in the literature [16, 17, 18].
The paper is organized as follows. In Sec. 2, we review the direct-sum QFT in Minkowski space-time established in [15]. DQFT of Minkowski space-time gives us the most valuable inputs to understand BHs, basically static curved space-times asymptotically Minkowski. The most important aspect of DQFT is the explicit realization of $\mathcal{C}\mathcal{P}\mathcal{T}$  invariance at every step of its construction, right from the way we quantize a field itself. DQFT presents a (quantum) picture of a conformal diagram that exploits further our understanding of time as a parameter in quantum theory. It lays a new foundation for understanding QFTCS. In Sec. 3, we discuss the SBH solution and the discrete symmetries. Then we point to the fact that the SBH in spherical coordinates is not the structure of metric if one wants to quantize a field in the background because of the coordinate singularity at the horizon $r=2GM$. Thus, it is necessary to choose a good set of coordinates, such as Kruskal-Szekers (KS). However, the most crucial observation is that the KS coordinates replace the horizons with discrete transformations. With the status of time being a parameter in quantum theory, the discrete space-time reflections need to be paid special attention here. We shall also comment on the assumptions made by Hawking before quantizing the fields in SBH space-time and emphasize the need for careful revision. In Sec. 4, we apply the DQFT formulation for the SBH in quantizing a KG field where we see an emergence of QFTCS that respects the discrete symmetries of SBH space-time and gives us an entirely new understanding of Hawking radiation. We clearly show here that, quantum mechanically, the SBH in KS coordinates presents us a picture of SBH space-time where there cannot be a popularly assumed quantum entanglement between the interior and exterior Hawking quanta. This is because the concept of time and the interior and exterior of SBH are not the same; one cannot use the same quantum theory rules everywhere. Moreover, the SBH metric is known to be static for $r\gtrsim 2GM$ and cosmological $r\lesssim 2GM$ as found in [19]. We witness a new structure of Fock space for SBH QFTCS, which separates (by superselection rule) the gravitational sub-system (SBH) from the environment, i.e., everything outside the SBH. Then a question arises, whether the interior and exterior Hawking quanta are independent of each other. To answer this question, we must take into account how the motion of a particle in the exterior of the SBH can gravitationally affect its interior counterpart. This will be a problem to discuss in detail and derive quantum mechanical consequences for it in the DQFT framework. Then we present a new picture for understanding (quantum) BH space-time and its implications for unitarity and observer complementarity, where we elucidate how there cannot be a Page curve [5, 6] and firewalls in the BH DQFT formalism. We add some comments about the evaporating BH DQFT picture and open questions for future investigation. In Sec. 5, we demonstrate that the BHs cannot be thought of as ordinary quantum systems, which is against the so-called central dogma [20] at the same time with BH DQFT we show that the three statements are true i.e., (i) Hawking radiation is described by a pure state, (ii) Low energy effective theory is valid, (iii) No firewalls, a free falling observer sees nothing unusual. Thus, our picture goes against the conjectural claims by AMPS [20]. Finally, in Sec. 6, we discuss the implications of our QFTCS to the formulation of S-matrix in BH space-times and also for quantum gravity. In Sec. 7, we conclude with a brief summary of results and some comments on future explorations in (quantum) BH physics. In Appendix. A, we present DQFT in Rindler spacetime [21] and discuss the similarities with the BH case.
2 A review of $\mathcal{C}\mathcal{P}\mathcal{T}$ invariance and direct-sum quantum field theory (DQFT) in Minkowski space-time
The main idea of this section is to give a review of DQFT formulation in Minkowski space-time, which was developed in [15]. An immediate question that can emerge here is, what is the need to look for QFT in Minkowski in a new way (which is our DQFT)? One of the prominent worries of (quantum) BH physics is what happens to the $\mathcal{C}\mathcal{P}\mathcal{T}$  invariance? It is the question that is explicitly raised in seminal works of Hawking’s, and also Gerard ’tHooft [2, 3, 22, 23, 24, 25]. The question of $\mathcal{C}\mathcal{P}\mathcal{T}$  has to take the central stage in BH physics because it is a successful outcome of SQFT. It is important to note that the discrete symmetries have played a central role in our understanding of particle (anti-particle) physics [26]. If we wish to quantize a field in the BH space-time, it is vital to remember the role of discrete space-time transformations. Indeed, the key observables of SQFT are the scattering amplitudes, which are $\mathcal{C}\mathcal{P}\mathcal{T}$  invariant. This means,
$\mathcal{A}_{N\to M}\left(p_{n},\,-p_{m}\right)=\mathcal{A}_{M\to N}\left(-p_{n},\,p_{m}\right)\,,$ (2.1)
where $\mathcal{A}_{N\to M}\left(p_{n},\,-p_{m}\right)$, $\mathcal{A}_{M\to N}\left(-p_{n},\,p_{m}\right)$ is the scattering amplitudes of N particles scattering into M particles and vice versa. Here $p_{n},,p_{m}$ denotes the 4-momenta of the $N$ and $M$ particles with $n=1,\cdots,N$ and $m=1,\cdots,M$. In (2.1), the pre and post-scattering states are determined by the sign of zeroth momenta, positive or negative. The $\mathcal{C}\mathcal{P}\mathcal{T}$  operation swaps the pre-scattering and the post-scattering states, i.e., the $\mathcal{P}\mathcal{T}$ operation changes the sign of zero the momenta but keeps the sign of 3-momenta and $\mathcal{C}$ operation swaps the particles with anti-particles [26]. Thus, the $\mathcal{C}\mathcal{P}\mathcal{T}$ , operation together turns the pre-scattering particles to the post-scattering anti-particles and vice versa. The $\mathcal{C}\mathcal{P}\mathcal{T}$  invariance given by (2.1) is a non-perturbative fact, i.e., it doesn’t matter how many loops are involved in the scattering process [26]. This is the remarkable feature of QFT in Minkowski space-time which is often not paid attention by the quantum gravity research community. Thus, the legitimate question is indeed what is the status of (2.1) in curved space-time?.
Now, let us recall that in the SQFT, the notion of pre-scattering (also called in states) states and post-scattering states (also called out states) are established by those in the limit $t\to-\infty$ and $t\to+\infty$ respectively. The same is reflected in the usual definition of the S-matrix by
[…]
where $T$ denotes the time ordering and $H_{int}$ is the interaction Hamiltonian.
[…]
2.1 Direct-sum quantum Mechanics
Before we present DQFT, we review here the construction of direct-sum quantum mechanics [^2] In the standard quantum mechanics, a positive energy state is given by $|\psi_{S}\rangle_{t}=|\psi_{S}\rangle_{0}e^{-iEt}$ if we presume arrow of time $t:-\infty\to\infty$, alternatively we can also define positive energy state $|\psi_{S}\rangle_{t}=|\psi_{S}\rangle_{0}e^{iEt}$ with opposite arrow of time $t:\infty\to-\infty$. Direct-sum QM defines a positive energy state without reference to the arrow of time (QM), which is based on formulating a single particle Hilbert space in terms of superselection sectors [^3] Superselection sectors are Hilbert spaces whose direct-sum forms the total Hilbert space. According to this, states in the superselection sectors cannot form a superposition with each other [30]. This construction has been discussed further in [15, 31].
The direct-sum Shrödinger equation where the wave function itself respects the $\mathcal{P}\mathcal{T}$ symmetry of the Hamiltonian
$i\frac{\partial}{\partial t}\begin{pmatrix}|\psi_{I}\rangle&\\ |\psi_{II}\rangle\end{pmatrix}=\begin{pmatrix}\hat{H}&0\\ 0&-\hat{H}\end{pmatrix}\begin{pmatrix}|\psi_{I}\rangle&\\ |\psi_{II}\rangle\end{pmatrix}\,,$ (2.3)
where
$|\psi\rangle=\frac{|\psi_{I}\rangle\oplus|\psi_{II}\rangle}{\sqrt{2}}=\frac{1}{\sqrt{2}}\begin{pmatrix}|\psi_{I}\rangle&\\ |\psi_{II}\rangle\end{pmatrix}\,.$ (2.4)
And
$\hat{H}|\psi_{I}\rangle$ $=\hat{H}_{I}\left(\hat{x}_{1},\hat{p}_{1}\right)|\psi_{I}\rangle$ (2.5)
$\hat{H}|\psi_{II}\rangle$ $=\hat{H}_{I}\left(\hat{x}_{2},\hat{p}_{2}\right)|\psi_{II}\rangle$
where [^4] Furthermore, note that $\Big{[}\hat{x}_{1},\,\hat{x}_{2}\Big{]}=\Big{[}\hat{p}_{1},\,\hat{p}_{2}\Big{]}=\Big{[}\hat{x}_{1},\,\hat{p}_{2}\Big{]}=\Big{[}\hat{p}_{1},\,\hat{x}_{2}\Big{]}=0\,.$ (2.6)
$=i\hbar\quad\hat{p}_{1}=-i\hbar\frac{\partial}{\partial x_{1}}\quad x_{1}=x\gtrsim 0\,$ (2.7)
$[\hat{x}_{2},\,\hat{p}_{2}]$ $=-i\hbar\quad\hat{p}_{2}=i\hbar\frac{\partial}{\partial x_{2}}\quad x_{2}=x\lesssim 0$
The factor of $\sqrt{2}$ in (2.4) is a normalization to have the total probabilities add up to 1 as
[…]
This clearly shows the concept of direct-sum, which is totally different from the usual summation in quantum theory. In (2.4), state vector $|\psi\rangle$ is defined in a Hilbert space, which can be written as
$\mathcal{H}=\mathcal{H}_{I}\oplus\mathcal{H}_{II}\,,$ (2.9)
where $\mathcal{H}_{I}$ and $\mathcal{H}_{II}$ are the Hilbert spaces (superselection sectors) corresponding to the spatial regions I and II that are parity conjugates. Note that the direct-sum description of quantum theory is quite similar to Wick, Wightman, and Wigner’s proposal of superselection rule [32, 30, 15]. One must not confuse direct-sum $\oplus$ with the usual summation $+$. When we take the direct-sum of operators, they take the block diagonal form like (2.12) and a direct-sum of two wave vectors [33, 34, 35].
In the direct-sum QM, a single quantum state (positive energy) is represented by direct-sum of two components where one evolves forward at position $x$ according to the arrow of time $t:-\infty\to\infty$ and the other evolves backward in time at position $-x$ according to the arrow of time $t:\infty\to-\infty$
$|\psi\rangle_{t}=\begin{pmatrix}|\psi_{I}\rangle_{0}e^{-iEt}\\ |\psi_{II}\rangle_{0}e^{iEt}\end{pmatrix}$ (2.10)
where $E$ is the energy eigenvalue in the case of the time-independent Hamiltonian. The wavefunction is
$\psi(x)=\langle x|\psi\rangle=\begin{pmatrix}\langle x_{1}|&\langle x_{2}|\end{pmatrix}\begin{pmatrix}|\psi_{1}\rangle\\ |\psi_{2}\rangle\end{pmatrix}=\frac{1}{2}\psi(x_{1})e^{-iEt}\Theta(t)+\frac{1}{2}\psi(x_{2})e^{iEt}\Theta(-t)$ (2.11)
where $\Theta(s)$ is heaviside-function.
2.2 Quantization of Klein-Gordon Field in DQFT
Here is the physical meaning of the direct-sum scheme of quantization.
[…]
The Klein-Gordon (KG) field operator in the DQFT formulation represented as a direct-sum of two components as written below
$\hat{\phi}\left(x\right)$ $=\frac{1}{\sqrt{2}}\hat{\phi}_{I}\left(t,\,\textbf{x}\right)\oplus\frac{1}{\sqrt{2}}\hat{\phi}_{II}\left(-t,\,-\textbf{x}\right)$ (2.12)
$=\frac{1}{\sqrt{2}}\begin{pmatrix}\hat{\phi}_{I}&0\\ 0&\hat{\phi}_{II}\end{pmatrix}$
where
$\hat{\phi}_{I}\left(t,\,\textbf{x}\right)$ $=\int\frac{d^{3}\mathbf{k}}{\left(2\pi\right)^{3/2}}\frac{1}{\sqrt{2\omega_{\mathbf{k}}}}\Bigg{[}\hat{a}_{I\,\mathbf{k}}e^{ik\cdot x}+\hat{a}^{\dagger}_{I\,\textbf{k}}e^{-ik\cdot x}\Bigg{]}$ (2.13)
$\hat{\phi}_{II}\left(-t,\,-\textbf{x}\right)$ $=\int\frac{d^{3}\mathbf{k}}{\left(2\pi\right)^{3/2}}\frac{1}{\sqrt{2\omega_{\mathbf{k}}}}\Bigg{[}\hat{a}_{II\,\mathbf{k}}e^{-ik\cdot x}+\hat{a}^{\dagger}_{II\,\textbf{k}}e^{ik\cdot x}\Bigg{]}\,,$
which represent components of field $\hat{\phi}$ in the spatial regions I and II defined by Parity $\mathcal{P}$ operation. Here the operators $a_{I},\,a_{II}$ and $a_{I}^{\dagger},\,a_{II}^{\dagger}$ are the annihilation and creation operators respectively satisfying the canonical commutation relations
$=[\hat{a}_{II\,\textbf{k}},\,\hat{a}_{II\,\textbf{k}^{\prime}}^{\dagger}]=\delta^{(3)}\left(\textbf{k}-\textbf{k}^{\prime}\right)$ (2.14)
$[\hat{a}_{I\,\textbf{k}},\,\hat{a}_{II\,\textbf{k}^{\prime}}]$ $=[\hat{a}_{I\,\textbf{k}},\,\hat{a}_{II\,\textbf{k}^{\prime}}^{\dagger}]=[\hat{a}_{I\,\textbf{k}}^{\dagger},\,\hat{a}_{II\,\textbf{k}^{\prime}}^{\dagger}]=0\,.$
and we demand the mixed commutation relations i.e., the second line of (2.14) such that
$[\hat{\phi}_{I}\left(x\right),\,\hat{\phi}_{II}\left(-x^{\prime}\right)]=0\,,$ (2.15)
Furthermore, the commutation relations of the field operator components $\hat{\phi}_{I}$,  $\hat{\phi}_{II}$ and the corresponding conjugate momenta are defined by
$[\hat{\phi}_{I}\left(t,\textbf{x}\right),\,\hat{\pi}_{I}\left(t,\,\textbf{x}^{\prime}\right)]=i\delta\left(\textbf{x}-\textbf{x}^{\prime}\right),\quad[\hat{\phi}_{II}\left(-t,-\textbf{x}\right),\,\hat{\pi}_{II}\left(-t,\,-\textbf{x}^{\prime}\right)]=-i\delta\left(\textbf{x}-\textbf{x}^{\prime}\right),$ (2.16)
where
${\pi}_{I}\left(t,\,\textbf{x}\right)=\frac{\partial\mathcal{L}_{\rm KG}}{\partial\left(\partial_{t}\phi_{I}\right)},\quad{\pi}_{II}\left(t,\,\textbf{x}\right)=-\frac{\partial\mathcal{L}_{\rm KG}}{\partial\left(\partial_{t}\phi_{II}\right)}$ (2.17)
The Fock space vacuum in the DQFT is the direct-sum of two
$|0\rangle_{T}=|0\rangle_{I}\oplus|0\rangle_{II}=\begin{pmatrix}|0\rangle_{I}\\ |0\rangle_{II}\ \end{pmatrix}\,.$ (2.18)
defined by
$\hat{a}_{I\,\textbf{k}}|0\rangle_{I}=0,\quad\hat{a}_{II\,\textbf{k}}|0\rangle_{II}=0\,.$ (2.19)
The total Fock space is also direct-sum of the two Fock spaces as
$\mathcal{F}_{T}=\mathcal{F}_{I}\oplus\mathcal{F}_{II}\,,$ (2.20)
where
$\mathcal{F}_{I}\left(\mathcal{H}\right)$ $=\bigoplus_{i=0}^{\infty}\mathcal{H}_{I}^{i}=\mathcal{C}\oplus\mathcal{H}_{I1}\oplus\left(\mathcal{H}_{I1}\otimes\mathcal{H}_{I2}\right)\oplus\left(\mathcal{H}_{I1}\otimes\mathcal{H}_{I2}\otimes\mathcal{H}_{I3}\right)\oplus\cdots$ (2.21)
$\mathcal{F}_{II}\left(\mathcal{H}\right)$ $=\bigoplus_{i=0}^{\infty}\mathcal{H}_{II}^{i}=\mathcal{C}\oplus\mathcal{H}_{II1}\oplus\left(\mathcal{H}_{II1}\otimes\mathcal{H}_{II2}\right)\oplus\left(\mathcal{H}_{II1}\otimes\mathcal{H}_{II2}\otimes\mathcal{H}_{II3}\right)\oplus\cdots\,,$
where $\mathcal{H}_{In}$ and $\mathcal{H}_{IIn}$ are nth particle Hilbert spaces. When we act the KG field operator on the total vacuum we obtain
$\hat{\phi}|0\rangle_{T}=\frac{1}{\sqrt{2}}\begin{pmatrix}\hat{\phi}_{I}&0\\ 0&\hat{\phi}_{II}\end{pmatrix}\begin{pmatrix}|0\rangle_{I}\\ |0\rangle_{II}\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}\hat{\phi}_{I}|0\rangle_{I}\\ \hat{\phi}_{II}|0\rangle_{II}\end{pmatrix}$ (2.22)
[…]
$[\mathcal{O}_{I}\left(t,\,\textbf{x}\right),\,\mathcal{O}_{II}\left(-t^{\prime},-\textbf{x}^{\prime}\right)]=0\,.$ (2.23)
This is a new condition in addition to the usual demand of causality conditions that operators corresponding to space-like distances should vanish. Here in DQFT, it holds for both Fock space operators.
$\left[\mathcal{O}_{I}\left(x\right),\,\mathcal{O}_{I}(y)\right]$ $=0,\quad\left(x-y\right)^{2}>0\,$ (2.24)
$\left[\mathcal{O}_{II}\left(x\right),\,\mathcal{O}_{II}(y)\right]$ $=0,\quad\left(x-y\right)^{2}>0\,.$
Note that $\hat{\phi}_{I}$ and $\hat{\phi}_{II}$ never get mixed by any interaction because of the definition of direct-sum (2.12). For example,
$\mathcal{L}_{int}=-\frac{\lambda}{3}\hat{\phi^{3}}=-\frac{\lambda}{3}\begin{pmatrix}\hat{\phi}_{I}^{3}&0\\ 0&\hat{\phi}_{II}^{3}\end{pmatrix}$ (2.25)
In SQFT the way we define a positive energy state is by following the usual way of writing Schrödinger equation as
$i\frac{\partial|\Psi\rangle}{\partial t}=\hat{H}|\Psi\rangle$ (2.26)
Quantum theory is symmetric under the change of $+i$ to $-i$ everywhere (time-symmetric). The S-matrix also becomes the direct-sum with two parts
$S_{T}$ $=S_{I}\oplus S_{II}$ (2.27)
where
[…]
with $T_{1},\,T_{2}$ representing the time orderings attached to the respective Fock space arrow of time. The scattering amplitude of DQFT becomes
$\mathcal{A}=\frac{\mathcal{A}_{I}^{N\to M}\left(p_{a},\,-p_{b}\right)+\mathcal{A}_{II}^{N\to M}\left(-p_{a},\,p_{b}\right)}{2}\,,$ (2.29)
where $\mathcal{A}_{I}^{N\to M}\left(p_{a},\,-p_{b}\right),\,\mathcal{A}_{II}^{N\to M}\left(-p_{a},\,p_{b}\right)$ are the amplitudes with respect to the Fock space $\mathcal{F}_{I}$ and $\mathcal{F}_{II}$ respectively and they are equal for scattering in Minkowski space-time. According to DQFT the scattering looks like $N$ particles coming from $\Big{\{}t\to-\infty,\,\textbf{x}\to-\infty\Big{\}}$ and scatttering to M particles in the limit $\Big{\{}t\to\infty,\,\textbf{x}\to\infty\Big{\}}$ according to the Fock space $\mathcal{F}_{I}$. While, according to Fock space $\mathcal{F}_{II}$ the picture is $N$ particles coming from $\Big{\{}t\to\infty,\,\textbf{x}\to\infty\Big{\}}$ and scatttering to M particles in the limit $\Big{\{}t\to-\infty,\,-\textbf{x}\to\infty\Big{\}}$. An observer who is part of the quantum system only sees the pre-scattering and post-scattering states, there is no notion of time associated with him/her. So the (LABEL:Amp) is observable and since
$\mathcal{A}_{I}^{N\to M}\left(p_{a},\,-p_{b}\right)=\mathcal{A}_{II}^{N\to M}\left(-p_{a},\,p_{b}\right)$ (2.30)
we arrive at the same observables of SQFT. Therefore, DQFT is no different from SQFT observationally, because of the $\mathcal{C}\mathcal{P}\mathcal{T}$  invariance.
2.3 Conformal diagram of DQFT in Minkowski space-time
Minkowski space-time is written in terms of the compactified coordinates by [36]
$ds^{2}=\frac{1}{\xi(T,R)}\left(-dT^{2}+dR^{2}+\frac{1}{4}\sin^{2}(2R)d\Omega^{2}\right)\,,$ (2.31)
where $-\pi+R<T<\pi-R$ and $-\pi+T<R<\pi-T$, and $\xi(T,R)$ is some positive function and $d\Omega^{2}$ represents the line element of two sphere.
$T$ $=\arctan\left(t+r\right)+\arctan\left(t-r\right)$ (2.32)
$R$ $=\arctan\left(t+r\right)-\arctan\left(t-r\right)$
which gives the following Minkowski space-time metric
We can notice that, under time reflection operation $\mathcal{T}$, we get
$\mathcal{T}:t\to-t\implies T\to-T\,\quad R\to R\,.$ (2.33)
The Fig. 1 represents the DQFT description of Minkowski space-time where we divide the space by Parity and assign opposite time evolution regions to the quantum fields in the spatial regions.
Figure 1: This figure is the depiction of Minkowski space-time in the DQFT construction, and we call it a "(quantum) conformal diagram" as it does not represent classical space-time. The right and left triangles are $\mathcal{P}\mathcal{T}$ conjugates of each other (region I and II, respectively). The quantum field operator $\hat{\phi}$ in DQFT is the direct-sum of $\hat{\phi}_{I},\hat{\phi}_{II}$ defined on the right and left, respectively, which follows from (2.12). The orange arrows represent the time direction.
We have several lessons from DQFT of Minkowski space-time which we need to keep in mind for QFTCS
[…]
  • Direct-sum is a useful mathematical operation to represent the full Hilbert space of physical systems spanned by sub-physical regions with different arrows of time. On the other hand direct-product (which is used to define entanglement in QM) is a mathematical operation of composite systems where the arrow of time is the same in both sub-system and the environment.
3 Schwarzschild Black Hole, coordinate systems and the discrete symmetries
The aim of this section is to discuss the SBH metric in different coordinates. Although this may be thought to be usual standard textbook material [36], our sincere appeal to the reader is that we specially pay attention to the discrete space-time symmetries of SBH space-time, which are pretty important for the understanding of quantum fields in such space-time. Therefore, every small detail of this section is extremely important for the latter understanding and formulation of QFTCS.
The SBH metric in the Schwarzschild coordinates can be written as
$ds^{2}=-\left(1-\frac{2GM}{r}\right)dt^{2}+\frac{1}{\left(1-\frac{2GM}{r}\right)}dr^{2}+r^{2}d\Omega^{2}\,,$ (3.1)
where $r>0$ is the radial coordinate $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\varphi^{2}$. We can clearly notice that (3.1) is static and it has the following discrete symmetries,
$\mathcal{T}:t\to-t,\quad\mathcal{P}:\left(\theta,\,\phi\right)\to\left(\pi-\theta,\,\pi+\varphi\right)$ (3.2)
which correspond to time reversal and parity operations. Notably, the $\mathcal{P}\mathcal{T}$ transformations (3.2) remains the same in the asymptotic Minkowski limit of the SBH metric (3.1) i.e., $r\to\infty$, therefore, we retain the definition (3.1) to be the crux for formulating quantum fields in SBH background.
But the form of metric (3.1) is not suitable everywhere to define quantum fields. It is because of the apparent coordinate singularity at $r=2GM$ which is certainly not a physical choice. Coordinate singularities are those where the curvature invariants are regular [^5] For example, the Kretschman scalar $\mathcal{K}=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{48G^{2}M^{2}}{r^{6}}$ for (3.1). , thus they are not the physical ones. An appropriate choice of coordinates is Kruskal-Szekers coordinates which are defined as the following. We request the reader to have careful attention to the following steps as there are not exactly what is understood from the textbook’s point of view.
First of all, one may think that since (3.1) is spherically symmetric, common wisdom is to ignore angular coordinates $\left(\theta,\,\phi\right)$ which we strongly request our reader not to do so. With this caution, let us define the so-called light cone coordinates $\left(u,\,v\right)$ as
$u=t-r_{\ast},\quad v=t+r_{\ast}\,,$ (3.3)
where $r_{\ast}$ is the so-called tortoise coordinate
$r_{\ast}=\frac{1}{2}\left(v-u\right)=r+2GM\ln\Bigg{|}\frac{r}{2GM}-1\Bigg{|},\,$ (3.4)
Here as
$r\to\infty\implies r_{\ast}\to\infty,\quad r\to 2GM\implies r_{\ast}\to-\infty,\quad r\ll 2GM\implies r_{\ast}\to r\,.$ (3.5)
The metric (3.1) now becomes
$ds^{2}=-\left(1-\frac{2GM}{r}\right)dudv+r^{2}d\Omega^{2}\,.$ (3.6)
The state of discrete symmetries (3.2) in the light cone coordinates become
$\mathcal{T}:t\to-t\implies u\to-v,\,v\to-u,\quad\mathcal{P}:\left(\theta,\,\phi\right)\to\left(\pi-\theta,\,\pi+\varphi\right)$ (3.7)
We can easily notice that the above transformation is very much similar to the one of Minkowski space-time in light cone coordinates discussed in [15].
The metric (3.6) is still not yet suitable to describe quantum fields because of the coordinate singularity again at $r=2GM$. It requires a non-linear coordinate transformation to obtain a form of metric that removes the apparent singularity at $r=2GM$. Let us first define
$U=-\frac{1}{4GM}e^{-\frac{u}{4GM}},\quad V=\frac{1}{4GM}e^{\frac{v}{4GM}}$ (3.8)
which are known as Kruskal-Szekers (KS) coordinates in which the metric (3.6) becomes
$ds^{2}=-\frac{2GM}{r}e^{1-\frac{r}{2GM}}dUdV+r^{2}d\Omega^{2}\,,$ (3.9)
where
$16G^{2}M^{2}e^{\frac{r}{2GM}}\left(1-\frac{r}{2GM}\right)=UV$ (3.10)
Using (3.10) we can express the radial coordinate as
$r=2GM\left(1+W\left(-\frac{UV}{16G^{2}M^{2}}\right)\right)\,,$ (3.11)
where $W$ is the Lambert function which satisfies $W\left(x\right)e^{W\left(x\right)}=x$.
According (3.10) we have
$r>2GM\implies UV<0,\quad r<2GM\implies UV>0\,.$ (3.12)
which shows a clear distinction between BH exterior and interior. Moreover, the KS coordinates are those that allow us to define the SBH metric (3.1) to the interior [36, 37]
Let us recall again (3.8) according to which $U<0,\,V>0$. Following (3.12) we can rather say
$r>2GM\implies\begin{cases}U=-\frac{1}{4GM}e^{-\frac{u}{4GM}}<0&V=\frac{1}{4GM}e^{\frac{v}{4GM}}>0\\ U=\frac{1}{4GM}e^{-\frac{u}{4GM}}>0&V=-\frac{1}{4GM}e^{\frac{v}{4GM}}<0\,.\end{cases}$ (3.13)
Examining carefully the first and second line of (3.13) we can see that the definition of KS coordinates changes by discrete transformation
$U\to-U\quad V\to-V\,$ (3.14)
The meaning of (3.14) is deeply associate with $\mathcal{P}\mathcal{T}$ transformations that we defined in (3.2). Subsequently, following (3.7) we can deduce that
$\mathcal{T}:$$\Bigg{\{}u\to-v,\quad v\to-u\Bigg{\}}\implies\Bigg{\{}U\to-U\quad V\to-V\Bigg{\}}\,,$ (3.15)
$\mathcal{P}:$ $\left(\theta,\,\phi\right)\to\left(\pi-\theta,\,\pi+\varphi\right)\,,$
which involves a redefinition that a function of $u$ should always be called $U(u)$ and the function of $v$ should always be called $V(v)$. Thus, the discrete transformation (3.14) is the $\mathcal{P}\mathcal{T}$ operation. Since it does not change anything in the metric (3.9), and since there are discrete space-time transformations involved, we should not at all think about them in terms of classical physics. In quantum theory, $\mathcal{P}\mathcal{T}$ transformations (along with $\mathcal{C}$) play a role in defining quantum fields which we recall again here. Similarly to (3.13) we can equally understand the BH interior with
$r<2GM\implies\begin{cases}U=\frac{1}{4GM}e^{-\frac{u}{4GM}}>0&V=\frac{1}{4GM}e^{\frac{v}{4GM}}>0\\ U=-\frac{1}{4GM}e^{-\frac{u}{4GM}}<0&V=-\frac{1}{4GM}e^{\frac{v}{4GM}}<0\,.\end{cases}$ (3.16)
which again presents two realizations of the interior with discrete flip of the KS coordinates (3.14). We can notice here that all different definitions of $\left(U,\,V\right)$ (3.13) and (3.16) are the whole symmetries of (3.9). Furthermore, (3.13) and (3.16) are associated with discrete flip of KS coordinates, one cannot get from one to another by any continuous space-time transformations.
Now we aim to understand the nature of SBH space-time (3.1). One of its peculiarities is that the points of
$\left(\theta,\,\phi\right)\Bigg{|}_{r=2GM},\,\left(\pi-\theta,\,\pi+\varphi\right)\Bigg{|}_{r=2GM}$ (3.17)
are space-like separated (see Fig. (2)) and it takes an infinite amount of energy for anything to cross $r=0$ and reach the other point [25]. This explains why one must be careful in suppressing the angular coordinates in SBH metric when one tries to make sense (quantum mechanically) of space-time.
Figure 2: This picture depicts how the SBH (3.1) rip apart the parity conjugate points on the sphere of radius $r=2GM$ to space-like distances.
Finally, we note that the meaning of near-horizon approximation
$r\to 2GM\implies\Big{\{}u\to\infty,\,v\to-\infty\Big{\}}\implies\begin{cases}U\to 0^{\mp}&V\to 0^{\pm}\,\text{for}\quad r>2GM\\ U\to 0^{\pm}&V\to 0^{\pm}\,\text{for}\quad r<2GM\end{cases}$ (3.18)
following (3.13) and (3.16).
3.1 An essential need to revise Hawking calculation
The idea of this sub-section is to highlight the importance of revising Hawking’s original calculation of 1975 [1]. Putting aside the remarkable result of Hawking’s paper, the need to save unitarity is important, and, in fact, Hawking himself kept working on it for several decades [22, 23]. Hawking pointed out, based on the success of AdS/CFT correspondence, that there should be a unitary evolution in (quantum) BH physics and he investigated the subject in the context of BMS symmetries in collaboration with Perry and Strominger [38, 39]. In the later Hawking works, the need for a careful analysis of the subject he initiated in 1975 with his seminal paper is suggested.
As we discussed previously, in this section, there are discrete symmetries of SBH, and one must pay careful attention from a quantum mechanical point of view. We note that in the context of QFTCS, quantum gravity, there can be a concept of time with no classical analog as it is explicitly discussed in the book by Carlo Rovelli [40]. One indication for this would be that the famous Wheeler-De Witt (WDW) equation, the most basic equation in the scope of (quantum) GR, is grappled with the problem of time because the time variable does not explicitly appear in the equation [41, 42, 43]. Therefore, the unitarity problem arising from the Hawking calculation signals a caveat to us that we are missing notice of an unforeseen feature related to the concept of time in the basic assumptions that are made at the beginning of understanding (quantum) BHs. This is precisely what this paper is about. In the previous section, we noticed several definitions of KS coordinates (3.13) and (3.16) that would leave the metric (3.9) invariant. In the original Hawking calculations, only the first line of (3.13) was taken into account and regarded as the BH formed by the gravitational collapse (the most popular classical Penrose diagram of it is depicted in Fig. (3)). This is the most important primary caveat in Hawking’s calculation because the classical and intuitive choice of coordinates and their values are put forward before quantization, regardless of the fact that "time" is a parameter in quantum theory. We know and understand so far the gravitational collapse from macroscopic, especially from GR point of view, and also some insights from microscopic physics in the sense at which point gravity would dominate from other fundamental forces of nature. But the whole process of collapse, formation of singularity (or formation of an object that closely resembles the BHs of GR [44, 45]) requires a clear understanding of scattering of fundamental degrees of freedom in the dynamical space-time. It is because of this reason that several investigations have been studying the high energy scattering during the collapse, especially quantum gravity frameworks ask the question of ultra-high energy scattering of strings and formation of BHs [46, 47, 48, 49, 50, 51]. Thus, we perhaps require a full understanding of quantum gravity before we can certainly answer what microscopic physics is involved in the formation of BHs. But the question of unitarity, in the context of quantum fields in curved space-time, takes precedence to the complete renormalizable theory of quantum gravity [13].
Having said all these, we return to the (classical) assumption made by Hawking in his work [1], which is restricting to a very particular definition of KS coordinates (3.8) and cutting by hand all the discrete symmetries of the KS metric (3.9) which are (3.13) and (3.16). This result of making a classical assumption of space-time and fixing it as depicted in the (classical) Penrose diagram of BH formed by dynamical collapse (See Fig. 3) is the origin of the unitarity problem. We have been institutionalized to think that time reversal in gravitational physics reverses whole dynamics, such as expanding the Universe to contracting the Universe and BH to the so-called White hole [36]. However, note that the first one needs to separate the concept of time classically from quantum mechanically. The classical understanding of space-time only instructs us whether there is an event horizon or an apparent horizon (i.e., a horizon changing with time). In recent works, it has been shown how quantum fields can evolve backward in time in an expanding Universe [15, 31, 52, 53]. In the next section, we describe how quantum fields in SBH can respect its full discrete symmetries of KS metric (3.9) and how unitarity emerges from there naturally. This follows from how we have constructed the DQFT of Minkowski space-time, which is explicitly $\mathcal{C}\mathcal{P}\mathcal{T}$  symmetric from its construction of quantizing fields.
Figure 3: The figure depicts the (classical) Penrose diagram of BH formed by gravitational collapse. Given that the concept of time in quantum physics is entirely different from the classical concept (according to GR), making an assumption on "time" before quantization is the origin of the unitarity problem in BH physics.
4 Application of DQFT for Schwarzschild Black Hole
This section uplifts the DQFT formalism to the context of SBH space-time in KS metric which is the suitable metric (3.9) for quantization. The idea of DQFT based on the discrete space-time transformations $\mathcal{P}\mathcal{T}$ makes use of the full description of the metric (3.9) with all realizations in (3.13) and (3.16). The physical interpretation of this emerges accordingly, which we shall discuss in detail in the later parts of the section. Here the BH with the horizon at $r=2M$ can be regarded as a gravitational subsystem and thus we expand the KG field in terms of two sets of creation and annihilation operators corresponding to the interior and exterior of SBH (still restricting the approximation $r\approx 2GM$). To be more clear about this, let us start with a KG in the SBH space-time
$S_{KG}=\int\sqrt{-g}d^{4}x\phi\square\phi\,.$ (4.1)
Since the SBH space-time (3.9) is spherically symmetric, it is useful to express the KG scalar field in terms of Spherical harmonics
$\phi\left(U,\,V,\,\theta,\phi\right)=\frac{1}{\sqrt{2}}\sum_{\ell m}\frac{\phi_{\ell m}}{r}Y_{\ell m}\left(\theta,\,\varphi\right)$ (4.2)
Computing the d’Alembertian for the metric (3.9) and integrating out on the $S_{2}$ sphere (4.1) can be rewritten as
[…]
where we used the orthogonality of spherical harmonics
$\int_{0}^{2\pi}\int_{0}^{\pi}Y_{\ell m}Y_{\ell^{\prime}m^{\prime}}\sin\theta d\theta d\varphi=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}\,.$ (4.4)
and $A(r)=\frac{2GM}{r}e^{1-\frac{r}{2GM}}$.
Note that our interest lies in the near horizon limits (3.18), which means
$\frac{1}{r}\frac{\partial r}{\partial U\partial V}$ $=\frac{1}{UV}\frac{W\left(-UV\right)}{\left(1+W\left(-UV\right)\right)^{4}}$ (4.5)
$\approx-\frac{1}{8G^{2}M^{2}}-\frac{5}{32G^{2}M^{2}}UV\cdots\,.$
which we can neglect together with the effective mass term in (4.3) for the BHs whose mass is significantly heavier than the Planck mass $\sim 10^{-5}g$. Then, we are left with an effectively massless scalar field in 1+1 dimensions. Note that by using near horizon approximation and expanding the scalar field as (4.2), we obtained an effective scalar $\phi_{\ell m}$ which propagates effectively in two-dimensional space-time. We may be tempted to see it as local Minkowski space-time, but we must distinguish it from the asymptotic Minkowski space-time of (3.1) in the limit $r\to\infty$.
To quantize the scalar $\phi_{\ell m}(U,\,V)$ which is a field component of the original field $\phi$ (See (4.2)) we need to promote it as an operator. Since we are quantizing a relativistic spherical wave (4.2), in the SBH space-time (3.9), we must not think about it in terms of (localized) particles, a common practice in the context of SQFT of Minkowski space-time. Nevertheless, the concept of a particle in curved space-time is a subject of separate investigation [54].
It is well-known from the classical understanding of SBH that the concept of time is totally different inside and outside the Schwarzschild radius $r_{S}=2GM$, in the sense that light cones of a classical point-like object would get rotated by 90°at $r_{S}=2GM$. Also, it has been shown that the exterior Schwarzchild metric is static, whereas the interior is cosmological in nature [19]. Before we do any quantization, we must take every observation and lesson from classical physics. Then, when we do quantization, we must respect aspects of discrete symmetries and leave aside the intuitions borrowed from classical physics. Since the nature of the metric in the BH interior and exterior is different (the concept of time is different), we must apply the superselection rule when we quantize the field. Now, what we do is to separate the spatial region of SBH (3.9) by the horizon
$\hat{\Phi}=\hat{\Phi}_{int}\oplus\hat{\Phi}_{ext}\,,$ (4.6)
where
$\Phi=\hat{\phi}_{\ell m},\quad\hat{\Phi}_{int}=\hat{\phi}_{\ell m}\Bigg{|}_{r\lesssim 2GM},\quad\hat{\Phi}_{ext}=\hat{\phi}_{\ell m}\Bigg{|}_{r\gtrsim 2GM}\,,$ (4.7)
are the quantum fields defined for the interior and exterior (but near horizon) of the SBH (3.1) represented by the KS metric (3.9). The KS metric has (discrete) symmetries that are intrinsic to quantum theory (once more, we recall that time is a parameter in quantum theory). The exterior of the SBH (3.1) is given by (3.13), while the interior of the SBH (3.1) is described by (3.16). Both in the interior and exterior, we have the discrete symmetries (3.15), which we use to formulate a direct-sum QFT for the interior and exterior regions of SBH. Therefore, we associate the above field operators to the two Fock spaces that describe the interior and exteriors of SBH. The direct-sum of these Fock spaces shall "bridge" the interior and the exterior.
$\mathcal{F}_{BH}=\mathcal{F}_{int}\oplus\mathcal{F}_{ext}\,.$ (4.8)
This is exactly the place where we differ from Hawking’s construction which uses a normal summation, i.e., $\hat{\phi}_{\ell m}=\hat{\phi}_{\ell m}^{ext}+\hat{\phi}_{\ell m}^{int}$ See. Eq. (2.7) of Hawking paper [3]. The simple change of normal summation to direct-sum operation has profound implications in resolving the problem of unitarity, which we shall discuss in Sec. 4.2.
Quantum theory is built on defining commutation and non-commutation relations, for example, (2.15), (2.16), (2.23) and (2.24) in the context of Minkowski DQFT. The most important lesson from QFT construction in Minkowski space-time is that the causality condition, which means the operators corresponding to spacelike distances, must vanish. In the context of SBH, we need to impose the operators corresponding to the space-like separated points (which are Parity conjugates. See Fig. 2) must commute. We shall show how one can achieve this in a while.
Then another question that arises here is what should be the commutator of
$\Big{[}\hat{\Phi}_{int},\,\hat{\Phi}_{ext}\Big{]}=?\,.$ (4.9)
According to Hawking’s conjecture, the above commutator is zero (See Eq. (2.8) of [3]). So Hawking assumes that the quantum fields created in the SBH interior (just inside the horizon) are totally independent of what is created in the exterior (just outside the horizon), and it is exactly the origin of the information problem. This has truly emerged from the "classical" intuition of BH no hair theorem according to classical GR [55, 56]. If the information paradox is to be solved or at least the first solution towards it has to emerge from the non-zero commutation of (4.9) which can be determined by a rather remarkable calculation of Dray and ’tHooft on the gravitational backreaction of (classical) point particles in GR [57]. We discuss this point further, based on the recent calculations of ’tHooft [58, 25, 59, 60, 61, 62, 63, 64, 65], which will help to determine (4.9).
4.1 Gravitational backreaction and quantum algebra
QFT has emerged with a clear understanding of classical special relativity and quantum mechanics. Both are brought together by a simple demand of commuting operators for space-like distances (2.24). Similarly, when we speak about quantum states created at $r\lesssim 2GM$ and $r\gtrsim 2GM$ in the context of BH evaporation [^6] It is well-known that Hawking quanta is the result of particle pair creation at $r\lesssim 2GM$ and $r\gtrsim 2GM$ and it is the effect of so-called stretched horizon as explained in detail by Mathur [16]. [16, 17, 18], it is essential to know from GR that if a (classical) particle at $r\lesssim 2GM$ can gravitationally affect its counterpart at $r\gtrsim 2GM$. The answer is ’Yes’ and is called the gravitational backreaction effect due to the shock waves created by (classical) particles in motion at the exterior and interior of the SBH (near horizon). This is the most famous calculation by Dray and ’tHooft [57] (See also works by Lousto and Sanchez [66]), which is well-summarized in Sec. 2 of [67]. The calculation of Dray and ’tHooft has been the central subject of investigation over the years in the context of gravitational memory effects, ultra-Planckian energy scatterings, gravitational collapse, and ultimately the formation of BHs [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 39, 78]. Despite the importance of Dray-’tHooft’ s paper highlighted in these later investigations about BHs, Dray-’tHooft´s result actually modifies the way we understand (quantum) BHs [^7] In this context ’tHooft in recent years emphasized the role of gravitational backreaction in understanding Hawking radiation [59, 63] and presented his new interpretation of (quantum) BH with antipodal identification [79] (i.e., classical topological change [25]) and later the reformed proposal with quantum clones [63]. We do not endorse these interpretations because they involve quite unusual boundary conditions and create problems for understanding the collapse and evaporation of BHs. Furthermore, the ’tHooft proposal completely removes the discussion about SBH interior and claims it is unimportant. However, we argue that the interior region of SBH is very important, and it plays a non-trivial role in the way we understand unitary BH radiation. The importance of BH interior from a quantum mechanical point of view is shown in recent works by Norma G. Sanchez [80, 81, 82] . The remarkable efforts of ’tHooft on the subject of gravitational backreaction and the corresponding quantum algebra [58, 25, 59, 60, 61, 83, 63, 65, 64] together with our DQFT formalism give us the more complete and concrete picture of (quantum) BHs. Below we shall discuss briefly the result of Dray-’tHooft paper [57] and also a few relevant parts of recent ’tHooft’ s remarkable non-commutative algebra associated with positions and momenta of the quantum mechanical particles in the interior and exterior of SBH. In this paper, we explain this within our context and notations to avoid confusion.
Note that we are interested in the near horizon limits given by (3.18) and how a (classical) massless particle at $r\lesssim 2GM$ can gravitationally affect the (classical) massless particle at $r\gtrsim 2GM$. This is the meaning of gravitational backreaction. Understanding this would tell us about the RHS of (4.9). Below, we apply the results of gravitational backreaction and the corresponding quantum algebra to our context [57, 67, 48, 59].
Suppose we have a massless (classical) point particle with momentum $P_{i}$ at an angle $\Omega$ moving in the direction of V near the horizon $r\lesssim 2GM$, it creates a shockwave $f\left(\Omega,\,\Omega^{\prime}\right)$ in the transverse direction of its motion described by [57, 67, 48]
$ds^{2}\approx-dUdV+\delta(U)f\left(\Omega,\,\Omega^{\prime}\right)dU^{2}+\left(2GM\right)^{2}d\Omega^{2}$ (4.10)
The delta function in the metric $\delta(U)$ signifies the location of the source at $U\approx 0$ corresponding to the energy-momentum tensor
$T_{UU}=P_{i}\delta(U)\delta(\Omega)$ (4.11)
The function $f(\Omega,\,\Omega^{\prime})$ is determined by solving the GR equations of motion in the near horizon approximation, and it gives
$(\Delta_{S}-1)f(\Omega,\,\Omega^{\prime})=-16\pi Gr_{S}^{2}e^{-1}P_{i}\delta^{2}\left(\Omega,\,\Omega^{\prime}\right)$ (4.12)
where $\Delta_{S}$ is the sphere Laplacian, $\delta^{2}\left(\Omega,\,\Omega^{\prime}\right)$, $\Omega\equiv\left(\theta,\,\varphi\right),\,\Omega^{\prime}\equiv\left(\theta^{\prime},\,\varphi^{\prime}\right)$ is the delta function and $r_{S}=2GM$. It was found that the massless (classical) point particle at the position $U_{i}$ (at an angle $\Omega^{\prime}$) near the horizon ($r\lesssim 2GM$) experience the shockwave (4.10) (backreaction) and consequently, its position gets shifted by an amount given by
$U_{e}\to U_{e}+8\pi Gr_{S}^{2}P_{i}f\left(\Omega,\,\Omega^{\prime}\right)$ (4.13)
Now if we perform a similar exercise with a massless particle with momentum $P_{e}$ moving along the direction U near the horizon ($r\lesssim 2GM$) at an angle $\Omega$, then it creates its own gravitational field with a shock wave, it causes a shift in the position of the massless (classical) point particle at the position $V_{i}$ near the horizon ($r\gtrsim 2GM$)
$V_{i}\to V_{i}+8\pi GP_{e}r_{S}^{2}f\left(\Omega,\,\Omega^{\prime}\right)\,.$ (4.14)
Since we are speaking about the near horizon approximation (3.18), it was found [59] that we can map the position of the massless point particle $V_{e}$ with the momentum of the massless (classical) particle $P_{i}$ and similarly map $U_{i}$ with $P_{e}$ as
$V_{i}=-\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}P_{e},\quad U_{e}=\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}P_{i}\,.$ (4.15)
The above mapping involves the expansion of positions and momenta with the spherical harmonics as
$V_{e}$ $\to\sum_{\ell m}V_{e}Y_{\ell m}\left(\theta,\,\phi\right),\quad U_{i}\to\sum_{\ell m}U_{i}Y_{\ell m}\left(\theta,\,\phi\right)$ (4.16)
$P_{e}$ $\to\sum_{\ell m}P_{e}Y_{\ell m}\left(\theta,\,\phi\right),\quad P_{i}\to\sum_{\ell m}P_{i}Y_{\ell m}\left(\theta,\,\phi\right)$
followed by appropriate dimensional rescaling (since $V_{e},U_{i}$ were dimensionless and $P_{e},\,P_{i}$ were momentum densities)
$V_{i}$ $\to r_{S}V_{i},\quad U_{e}\to r_{S}U_{e}\,$ (4.17)
$P_{e}$ $\to r_{S}^{-3}P_{e},\quad P_{i}\to r_{S}^{-3}P_{i}\,.$
Note that in the above steps, we ignore all the effects of transverse momenta since for sufficiently large spherically symmetric SBH the transverse effects can be negligible [63, 65].
Given now that GR clearly teaches us how a (classical) massless particle moving along the direction of V can gravitational affect the particle moving along the direction of U through gravitational backreaction due to shock waves (it is also known as the Shapiro effect), ’tHooft has applied the rules of quantum mechanics to find the following commutation relations which form the basis of BH QFT construction and ultimately allow us to find an answer to our question (4.9).
$\Big{[}\hat{V}_{i},\,\hat{P}_{i}\Big{]}=i\hbar,\quad\Big{[}\hat{U}_{e},\,\hat{P}_{e}\Big{]}=i\hbar$ (4.18)
As a consequence of (4.15) we end up with a most curious relation that forms the basis for the combined effect of gravity and quantum mechanics
$\Big{[}\hat{V}_{i},\,\hat{U}_{e}\Big{]}=i\hbar\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)},\quad\Big{[}\hat{P}_{i},\,\hat{P}_{e}\Big{]}=-i\hbar\frac{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}{8\pi G}$ (4.19)
Recalling that QFT is nevertheless an extension of quantum mechanics with relativistic concepts [26], we can already read the RHS of (4.9) from the RHS of (4.19). The rather very curious thing about the relations (4.19) is that it involves a combination of $G,\,\hbar,\,c$ and also Schwarzschild radius $r_{S}$, which signals that BHs are the key to understanding quantum gravity [83].
4.2 BH DQFT, unitary Hawking radiation, and conformal diagram
In (4.6) and (4.8), we have the field operator and the respective Fock space written as a direct-sum of the two components corresponding to the exterior and interior of the SBH (3.1) i.e.,  $r\gtrsim 2GM$ and $r\lesssim 2GM$. Now, in a similar way to how we have DQFT of Minkowski space-time, based on the discrete symmetries of the Minkowski metric (Sec. 2), we can formulate BH DQFT based on the discrete space-time transformations (3.15) both for the interior and exterior of SBH as follows.
In the view of (3.15), (3.13) and (3.16) and BH DQFT, we expand (4.6) as
$\hat{\Phi}$ $=\hat{\Phi}_{int}\oplus\hat{\Phi}_{ext}$ (4.20)
$=\frac{1}{\sqrt{2}}\left(\hat{\Phi}^{I}_{int}\oplus\hat{\Phi}^{II}_{int}\right)\oplus\frac{1}{\sqrt{2}}\left(\hat{\Phi}^{I}_{ext}\oplus\hat{\Phi}^{II}_{ext}\right)\,$
where
$\hat{\Phi}^{I}_{ext}$ $=\int_{-\infty}^{\infty}\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2|k|}}\Bigg{[}\hat{b}_{Ik}e^{-i|k|T+ikX}+\hat{b}_{Ik}^{\dagger}e^{i|k|T-ikX}\Bigg{]}$ (4.21)
$=\int_{0}^{\infty}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\Bigg{[}\hat{b}_{I\omega}e^{-i\omega U}+\hat{b}_{I-\omega}e^{-i\omega V}+\hat{b}_{I\omega}^{\dagger}e^{i\omega U}+\hat{b}_{I-\omega}^{\dagger}e^{i\omega V}\Bigg{]}$
$\hat{\Phi}^{II}_{ext}$ $=\int\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2|k|}}\Bigg{[}\hat{b}_{IIk}e^{i|k|T-ikX}+\hat{b}_{IIk}^{\dagger}e^{-i|k|T+ikX}\Bigg{]}$
$=\int_{0}^{\infty}\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\Bigg{[}\hat{b}_{II\omega}e^{i\omega U}+\hat{b}_{II-\omega}e^{i\omega V}+\hat{b}_{II\omega}^{\dagger}e^{-i\omega U}+\hat{b}_{II-\omega}^{\dagger}e^{-i\omega V}\Bigg{]}$
$\hat{\Phi}^{I}_{int}$ $=\int\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2|k|}}\Bigg{[}\hat{c}_{Ik}e^{i|k|T-ikX}+\hat{c}_{Ik}^{\dagger}e^{-i|k|T+ikX}\Bigg{]}$
$\hat{\Phi}^{II}_{int}$ $=\int\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2|k|}}\Bigg{[}\hat{c}_{IIk}e^{-i|k|T+ikX}+\hat{c}_{IIk}^{\dagger}e^{i|k|T-ikX}\Bigg{]}$
where $\omega=|k|$. Here
$T=\frac{U+V}{2},\quad X=\frac{V-U}{2}\,.$ (4.22)
The creation and annihilation operators above satisfy the following commutation relations
$\Big{[}\hat{b}_{I\,k},\,\hat{b}_{I\,k^{\prime}}^{\dagger}\Big{]}$ $=\Big{[}\hat{b}_{II\,k},\,\hat{b}_{II\,k^{\prime}}^{\dagger}\Big{]}=\hbar\delta^{(3)}\left(k-k^{\prime}\right)$ (4.23)
$\Big{[}\hat{b}_{I\,k},\,\hat{b}_{II\,k^{\prime}}\Big{]}$ $=\Big{[}\hat{b}_{I\,k},\,\hat{b}_{II\,k^{\prime}}^{\dagger}\Big{]}=\Big{[}\hat{b}_{I\,k}^{\dagger},\,\hat{b}_{II\,k^{\prime}}^{\dagger}\Big{]}=0\,.$
$\Big{[}\hat{c}_{I\,k},\,\hat{c}_{I\,k^{\prime}}^{\dagger}\Big{]}$ $=\Big{[}\hat{c}_{II\,k},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=\hbar\>\>\delta^{(3)}\left(k-k^{\prime}\right)$
$\Big{[}\hat{c}_{I\,k},\,\hat{c}_{II\,k^{\prime}}\Big{]}$ $=\Big{[}\hat{c}_{I\,k},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=\Big{[}\hat{c}_{I\,k}^{\dagger},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=0\,.$
Here the 2nd and 4th line of commutation relations (4.23) leads to
$\Big{[}\hat{\Phi}_{Iext},\,\hat{\Phi}_{IIext}\Big{]}=0$ (4.24)
$\Big{[}\hat{\Phi}_{Iint},\,\hat{\Phi}_{IIint}\Big{]}=0\,.$
which is similar to the one (2.23). This condition takes care of operators at space-like distances who should commute (See Fig. 2). As a consequence of (4.19) the exterior and interior quantum field operators should be non-commutative as
$\Big{[}\hat{\Phi}_{Iext},\,\hat{\Phi}_{Iint}\Big{]}$ $=i\hbar\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}$ (4.25)
$\Big{[}\hat{\Phi}_{IIext},\,\hat{\Phi}_{IIint}\Big{]}$ $=i\hbar\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}\,.$
and we also demand
$\Big{[}\hat{\Phi}_{Iext},\,\hat{\Phi}_{IIint}\Big{]}=0$ (4.26)
$\Big{[}\hat{\Phi}_{Iint},\,\hat{\Phi}_{IIext}\Big{]}=0\,.$
which is because the operators that correspond to opposite time evolutions and space-like distances (See Fig. 4 and Fig. 2) must vanish. So (4.25) and (4.26) imply
$\Big{[}\hat{b}_{I\,k},\,\hat{c}_{I\,k^{\prime}}\Big{]}$ $=\Big{[}\hat{b}_{II\,k}^{\dagger},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=i\hbar\frac{8\pi G}{r_{S}^{2}\left(\ell^{2}+\ell+1\right)}\delta^{(3)}\left(k-k^{\prime}\right)$ (4.27)
$\Big{[}\hat{b}_{I\,k},\,\hat{c}_{II\,k^{\prime}}\Big{]}$ $=\Big{[}\hat{b}_{I\,k},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=\Big{[}\hat{b}_{I\,k}^{\dagger},\,\hat{c}_{II\,k^{\prime}}^{\dagger}\Big{]}=0\,.$
In the near-horizon limit (3.18) the quantum vacuum of SBH as per the decomposition (4.21) is a direct-sum as
$|0\rangle_{BH}$ $=\Big{(}|0\rangle_{Iint}\oplus|0\rangle_{IIint}\Big{)}\oplus\Big{(}|0\rangle_{Iext}\oplus|0\rangle_{IIext}\Big{)}$ (4.28)
where the components of vacuum (4.28) defined by
$\hat{b}_{Ik}|0\rangle_{Iext}$ $=0,\quad\hat{b}_{IIk}|0\rangle_{IIext}=0$ (4.29)
$\hat{c}_{Ik}|0\rangle_{Iint}$ $=0,\quad\hat{c}_{IIk}|0\rangle_{IIint}=0\,.$
The total Fock space is again a direct-sum corresponding to the different spatial regions and arrows of time ($r\lesssim$ and $r\gtrsim 0$)
$\mathcal{F}_{BH}=\mathcal{F}_{Iint}\oplus\mathcal{F}_{IIint}\oplus\mathcal{F}_{Iext}\oplus\mathcal{F}_{IIext}$ (4.30)
Thus, there are 4 superselection sectors because there are 4 discrete transformations (3.13) and (3.16). In all the above expressions, the following meaning of notations is implicit
  • Region $I_{int}$ is the interior spatial region of SBH $r\lesssim 2GM$ with $U>0,\,V>0$ i.e., half of the interior. The vacuum in this region is $|0\rangle_{Iint}$.
  • Region $II_{int}$ is the interior spatial region of SBH $r\lesssim 2GM$ with $U>0,\,V>0$ i.e., remaining half of the interior. The vacuum in this region is $|0\rangle_{IIint}$.
  • Region $I_{ext}$ is the exterior spatial region of SBH $r\gtrsim 2GM$ with $U<0,\,V>0$ i.e., half of the exterior. The vacuum in this region is $|0\rangle_{Iext}$.
  • Region $II_{ext}$ is the exterior spatial region of SBH $r\gtrsim 2GM$ with $U>0,\,V<0$ i.e., remaining half of the exterior. The vacuum in this spatial region is $|0\rangle_{IIext}$.
All this means is in the exterior i.e., $r\gtrsim 2GM$, the quantum state
$\hat{\Phi}_{ext}|0\rangle_{ext}$ $=\frac{1}{\sqrt{2}}\left(\hat{\Phi}_{Iext}\oplus\hat{\Phi}_{IIext}\right)\Big{(}|0\rangle_{Iext}\oplus|0\rangle_{IIext}\Big{)}$ (4.31)
$=\frac{1}{\sqrt{2}}\left(\hat{\Phi}_{Iext}|0\rangle_{Iext}\oplus\>\hat{\Phi}_{IIext}|0\rangle_{IIext}\right)$
is described by two components $\hat{\Phi}_{Iext}|0\rangle_{Iext}$ and $\hat{\Phi}_{IIext}|0\rangle_{IIext}$. Similarly, in the interior, i.e, $r\lesssim 2GM$, the quantum state is described by two components
$\hat{\Phi}_{int}|0\rangle_{ext}$ $=\frac{1}{\sqrt{2}}\left(\hat{\Phi}_{Iint}\oplus\hat{\Phi}_{IIint}\right)\Big{(}|0\rangle_{Iint}\oplus|0\rangle_{IIint}\Big{)}$ (4.32)
$=\frac{1}{\sqrt{2}}\left(\hat{\Phi}_{Iint}|0\rangle_{Iint}\oplus\>\hat{\Phi}_{IIint}|0\rangle_{IIint}\right)$
We represent these quantum fields in the conformal diagram Fig. 4, which illustrates the SBH DQFT, very similar to how we represented DQFT of Minkowski space-time in Fig. 1.
For an observer at $r\to\infty$ ($r_{\ast}\to\infty$) of SBH space-time (3.1), the quantum field is in Minkowski space-time and in the DQFT formalism we have
$\hat{\Phi}_{ext}=\frac{1}{\sqrt{2}}\left(\hat{\Phi}_{I\infty}\oplus\hat{\Phi}_{II\infty}\right)\,,$ (4.33)
where
$\hat{\Phi}_{I\infty}$ $=\int_{0}^{\infty}\frac{d\tilde{\omega}}{\sqrt{2\pi}}\frac{1}{\sqrt{2\tilde{\omega}}}\Bigg{[}\hat{a}_{I\tilde{\omega}}e^{-i\tilde{\omega}u}+\hat{a}_{I-\tilde{\omega}}e^{-i\tilde{\omega}v}+\hat{a}_{I\tilde{\omega}}^{\dagger}e^{i\tilde{\omega}u}+\hat{a}_{I-\tilde{\omega}}^{\dagger}e^{i\tilde{\omega}v}\Bigg{]}$ (4.34)
$\hat{\Phi}_{II\infty}$ $=\int_{0}^{\infty}\frac{d\tilde{\omega}}{\sqrt{2\pi}}\frac{1}{\sqrt{2\tilde{\omega}}}\Bigg{[}\hat{a}_{II-\tilde{\omega}}e^{i\tilde{\omega}u}+\hat{a}_{II-\tilde{\omega}}e^{i\Omega v}+\hat{a}_{II\tilde{\omega}}^{\dagger}e^{-i\tilde{\omega}u}+\hat{a}_{II\tilde{\omega}}^{\dagger}e^{-i\tilde{\omega}v}\Bigg{]},$
where the creation and annihilation operators satisfy
$\Big{[}\hat{a}_{I\,\tilde{\omega}},\,\hat{a}_{I\,\tilde{\omega}^{\prime}}^{\dagger}\Big{]}$ $=\Big{[}\hat{a}_{II\,\tilde{\omega}},\,\hat{a}_{II\,\tilde{\omega}^{\prime}}^{\dagger}\Big{]}=\hbar\delta^{(3)}\left(\tilde{\omega}-\tilde{\omega}^{\prime}\right)$ (4.35)
$\Big{[}\hat{a}_{I\,\tilde{\omega}},\,\hat{a}_{II\,\tilde{\omega}^{\prime}}\Big{]}$ $=\Big{[}\hat{a}_{I\,\tilde{\omega}},\,\hat{a}_{II\,\tilde{\omega}^{\prime}}^{\dagger}\Big{]}=\Big{[}\hat{a}_{I\,\tilde{\omega}}^{\dagger},\,\hat{a}_{II\,\tilde{\omega}^{\prime}}^{\dagger}\Big{]}=0\,.$
This asymptotic Minkowski vacuum is given by
$|0\rangle_{\infty}=|0\rangle_{I\infty}\oplus|0\rangle_{II\infty},$ (4.36)
where
$\hat{a}_{I\tilde{\omega}}|0\rangle_{I}=0,\quad\hat{a}_{II\,\tilde{\omega}}|0\rangle_{II\infty}=0\,.$ (4.37)
Figure 4: This represents the conformal diagram of BH DQFT. Here $I_{ext}$ describes an exterior region of SBH spanned $\left(\theta,\,\varphi\right)$ where a component of the quantum field $\hat{\Phi}_{I\,ext}|0\rangle_{Iext}$ evolves forward in time while the $II_{ext}$ described the exterior region spanned by the parity conjugate points $\left(\pi-\theta,\,\pi+\varphi\right)$ where the second component $\hat{\Phi}_{II\,ext}|0\rangle_{IIext}$ of quantum field evolves backward in time. Similarly, $I_{int}$ describes interior region of SBH where $\hat{\Phi}_{I\,int}|0\rangle_{Iext}$ evolves forward in time while the $II_{int}$ describes the interior region where the second component of quantum field $\hat{\Phi}_{II\,int}|0\rangle_{IIint}$ evolves backward in time. The arrows represent the direction of the time evolution of a quantum field in each region. Every point in the yellow shared area represents the points $\left(\theta,\,\varphi\right)$, whereas the green shared regions represent the parity conjugated points $\left(\pi-\theta,\,\pi+\varphi\right)$. The red line represents $r=0$ singularity and is identified in the upper and lower half of the interior regions.
By applying the technique of Bogoliubov transformation, we obtain the number density of particles created by operators $\hat{a}^{\dagger}_{I\,k},\,\hat{a}^{\dagger}_{II\,k}$ in the vacuum $|0\rangle_{BH}$ which is what exactly asymptotic observer see as Hawking particles. Following [84] we can compute
[…]
where
$\alpha_{\omega\tilde{\omega}}$ $=\sqrt{\frac{\tilde{\omega}}{\omega}}\int_{0}^{\infty}\frac{du}{2\pi}e^{i\tilde{\omega}u-i\tilde{\omega}U}\quad\beta_{\omega\tilde{\omega}}$ $=\sqrt{\frac{\tilde{\omega}}{\omega}}\int_{0}^{\infty}\frac{du}{2\pi}e^{i\tilde{\omega}u+i\tilde{\omega}U}$ (4.39)
$\tilde{\alpha}_{\omega\tilde{\omega}}$ $=\sqrt{\frac{\tilde{\omega}}{\omega}}\int_{0}^{\infty}\frac{du}{2\pi}e^{-i\tilde{\omega}u+i\tilde{\omega}U}\quad\tilde{\beta}_{\omega\tilde{\omega}}$ $=\sqrt{\frac{\tilde{\omega}}{\omega}}\int_{0}^{\infty}\frac{du}{2\pi}e^{-i\tilde{\omega}u-i\tilde{\omega}U}$
Given all the above relations, the asymptotic observer sees the number density
$N_{\tilde{\omega}}$ $=\frac{1}{2}{}_{BH}\langle 0|\left(\hat{a}_{I\,\tilde{\omega}}^{\dagger}\oplus\hat{a}_{II\,\tilde{\omega}}^{\dagger}\right)\left(\hat{a}_{I\,\tilde{\omega}}\oplus\hat{a}_{II\,\tilde{\omega}}\right)|0\rangle_{BH}$ (4.40)
[…]
$=\frac{1}{e^{8\pi GM\omega/\hbar}-1}\,,$
where the prefactor $\frac{1}{2}$ is because of the prefactor in (4.33).
Applying thermodynamics the Hawking temperature of the BH is
$T=\frac{\hbar}{8\pi GM}$ (4.41)
Note that once we derive the number density of the particles that a distant observer can see, we can apply (classical) thermodynamics to understand how BH evaporates following the footsteps of Hawking and Bekenstein [1, 85]. Following the Stefan-Boltzmann law, the mass of the BH decreases as
$\frac{dM}{dt}=-\frac{n_{\gamma}}{15360\pi M^{2}}\,,$ (4.42)
where $n_{\gamma}$ denotes the number of degrees of freedom. Applying the second law of thermodynamics, the entropy of the BH is estimated to be [84]
$S_{BH}=4\pi M^{2}\,.$ (4.43)
In the next sections, we show that the Hawking radiation we derived here with DQFT gives a thermal spectrum with pure states. This is what is needed to achieve unitarity and observer complementarity. In a recent paper, we have studied DQFT in Rindler spacetime [21], and we have shown that one can achieve a thermal spectrum for every observer in Rindler spacetime preserving unitarity. Some of these details are briefly presented in Appendix. A. Especially it is interesting to notice similarities between Fig. 5 and 4 which elucidates universal features of DQFT in different spacetimes with horizons.
4.2.1 Unitarity and BH DQFT solution to the information paradox
From the beginning of understanding (quantum) BHs [1, 3], it has been assumed that the Hawking radiation is a pair production of particles in the interior (very near to the horizon $r\lesssim 2GM$) and the exterior (very near to the horizon $r\gtrsim 2GM$) and their combined Hilbert space is the direct product of individual Hilbert spaces [86, 16, 18]
$\mathcal{H}_{P}=\mathcal{H}_{int}\otimes\mathcal{H}_{ext}$ (4.44)
This originates an external particle that a distant observer sees in a mixed state. This has been projected to originate the unitarity violation, given the assumption of BH formation with pure states [5, 11, 9]. Furthermore, it was shown by Don Page that the entanglement entropy between the particle pair grows and can exceed the entropy of the SBH when half of the BH mass has evaporated [5, 6]. To save unitarity, Don Page proposed that the entanglement entropy of the Hawking radiation must start to decrease after some time (Page time) so that the unitarity is recovered after the complete evaporation of BH [5, 6]. This is called the Page curve [20] and to obtain Page curve, several quantum gravity proposals have emerged in recent years [11, 87].
From our point of view, the failure of unitarity, in the sense of Hawking radiation, is a mixed state, and the growth of entanglement entropy completely relies on the Hilbert space structure of Hawking pair (4.44). The BH DQFT proposal leads to the direct-sum-Fock space (4.8) and it leads to the Hilbert space of the Hawking pair as
$\mathcal{H}_{\mathcal{P}}$ $=\mathcal{H}_{int}\oplus\mathcal{H}_{ext}\,,$ (4.45)
This leads to a pure (maximally) entangled state becomes direct-sum of two pure states of the superselection sectors $\left(\mathcal{H}_{int},\,\mathcal{H}_{ext}\right)$
$|\tilde{\psi}_{12}\rangle=\sum_{m,n}\tilde{c}_{mn}|\tilde{\phi}_{m1}\rangle\otimes|\tilde{\phi}_{n2}\rangle=\frac{1}{\sqrt{2}}\sum_{m,n}\tilde{c}_{mn}\left(|\tilde{\phi}^{\rm ext}_{m1}\rangle\otimes\tilde{\phi}^{\rm ext}_{n2}\rangle\right)\oplus\left(|\tilde{\phi}^{\rm int}_{m1}\rangle\otimes\tilde{\phi}^{\rm int}_{n2}\rangle\right)$ (4.46)
where $\tilde{c}_{mn}\neq\tilde{c}_{n}\tilde{c}_{m}$, $|\phi_{1}\rangle=\sum_{m}\tilde{c}_{m}|\phi_{m1}\rangle$ and $|\phi_{2}\rangle=\sum_{n}\tilde{c}_{n}|\phi_{n2}\rangle$, the superscripts "ext" mean field component for $r\gtrsim 2GM$ whereas "int" mean field component for $r\lesssim 2GM$. This implies Hawking’s pair of quanta (pure state) is divided into two pure state components of the superselection sectors. The rule of direct-sum separates the Hilbert spaces into two separate spatial regions (superselection sectors), and we cannot have a superposition of states between those regions. This is very similar to the superselection rule proposed by Wick, Wightman, and Wigner [32]. Due to the Hilbert space structure (4.45), we do not have any Page curve and no unitarity violation. We get unitary physics for quantum states in the interior and exterior. A similar thing happens when one does DQFT in Rindler spacetime [21], which we briefly discuss in Appendix. A.
Coming to the information problem question basically arises from the fact that the creation and annihilation operators of the interior and exterior field operators are taken to commute (See Eq. (2.8)-(2.10) of [3]). But assuming ’tHooft’s quantum commutation relations, derived from gravitational backreaction (4.19), we arrived at (4.25) which relates interior and exterior fields in a very non-trivial way. This is the first step to solving the information problem. This means that by observing Hawking radiation from outside, we should be able to make computations to probe the interior. This subject needs further investigation of BH DQFT and we leave it for future investigations.
4.2.2 Black hole is not an ordinary quantum system
According to the so-called central dogma [20], BH must be an ordinary quantum system. Actually, this means the Hilbert space of the degrees of freedom of the subsystem and the environment must be entangled by direct-product operation (4.44). This means the concept of time should be the same in the subsystem and the environment. This is not the situation for BH, and we argue that BH cannot be an ordinary quantum system, and our proposal goes against the central dogma.
If the central dogma is not true, it was conjectured that we should see so-called firewalls [20] i.e., a freely falling observer should see particles that violate the equivalence principle. We ascertain that this is not our case because our near Horizon limit (3.18) vacuum (4.28) is nearly Minkowski for a local observer because it is completely $\mathcal{C}\mathcal{P}\mathcal{T}$  invariant [^8] Since we are quantizing a real scalar field, charge conjugation $\mathcal{C}$ does not play a role but it is easy to extend our formalism for complex scalar fields. (following (3.15)). Therefore, BH DQFT is a possible solution to have unitary quantum physics without firewalls and without Page time.
5 Questions of (quantum) black hole formation and evaporation
The idea of this section is to make a logical guess on what a conformal diagram of BH DQFT can look like. Classically, the description of the BH space-time diagram formed by collapse looks like Fig. 3 as Hawking initially described [1]. However, it is very unclear how to give a quantum mechanical (microscopic) description of collapse. During the collapse, we have matter imploding, and the horizon of the BH grows with time, which is expected to be a much more rapid process than evaporation when the horizon size starts to shrink. Of course, for the high energy modes, collapse can be described as an ultra-high energy superplanckian scattering [69, 49, 50, 51, 80, 81, 82]. It is still an open question to understand the transition between Minkowski space-time QFT and the context when gravitational effects on quantum fields cannot be neglected. To be precise, the scattering of particles under the effect of dynamical space-time is the key issue to be understood. Collapse or evaporation of a (quantum) BH means that, macroscopically, there is a direction of the system time evolution encoded in the growing or shrinking of the BH horizon. The macroscopic dynamical effects will break $\mathcal{C}\mathcal{P}\mathcal{T}$ symmetry for the microscopic degrees of freedom. A violation of time reversal symmetry should distort the (Not Penrose) diagram of BH, where the regions where the components of fields evolve backward in time behave differently from the components of fields that evolve forward in time. This is similar to the recent study of how quantum fields in an expanding Universe would create parity asymmetry in cosmic microwave background [31, 53, 52]. We anticipate a similar effect for BH physics, but we seriously lack any machinery to do such computations. We leave this for future investigations.
6 BH DQFT implications for S-matrix and quantum gravity
S-matrix is the most important notion in QFT, and it is supposed to describe the scattering of particles. The notion of S-matrix has varied meanings in curved space-time, and most often, it is not considered as a required concept. However, some questions remain, such as: how do the standard model particles we know in nature react to curved space-times with horizons? How does a particle, and if it decays, get affected by the presence of curved space-time? Therefore, if not the S-matrix, we need something similar to describe particle scatterings. There is another notion of S-matrix that has to do with the complete formation and full evaporation of BH, which definitely requires a full understanding of quantum gravity and currently, we do not have any definite candidate for that. Given we have multiple frameworks of quantum gravity [88, 10] and currently we essentially need quantitative ideas to test quantum gravity with astrophysical and gravitational wave observations to anchor what direction is correct for quantum gravity. Even to make significant advancements in quantum gravity research, we do need to address the issue of quantum fields in curved space-time.
BH DQFT quantization gives an elegant way to separate the BH interior from the exterior without violating unitarity and, at the same time, thanks to the ’tHooft quantum backreaction algebra, we do know now how interior and exterior degree of freedom are connected with each other non-trivially (4.25). With this, we can aim to understand first particle interactions, scattering, and decays near the horizon. To an exterior observer, all the particle interactions expected happen unitarily, of course, if degrees of freedom disappear in the interior of the BH they would definitely leave an imprint in the behavior of the degrees of freedom outside by (4.25) or its modification when we include interactions.
7 Conclusions and outlook
In this section, we aim to give a quick summary of what was done in this paper. Since the proposal of Hawking’s seminal paper on BHs, we have been highly grappled with issues related to information paradox, unitarity violation, and breakdown of predictability. All of these emerged from quantizing a scalar field in SBH space-time. The identification of this problem even dates back to 1935, when Einstein and Rosen pointed out the "Particle problem in General Relativity" [37], which exposes the fundamental problem between gravity and quantum mechanics. Several developments have happened in the last decades under the scope of quantum gravity to understand BHs and the several enigmas associated with them. Don Page’s calculation of the entanglement entropy growth beyond the BH entropy has bolstered even more the puzzles associated with BHs. In this paper, we paid very critical attention to the origin of inconsistencies in our understanding of BHs with quantum mechanics. We noticed the importance of $\mathcal{C}\mathcal{P}\mathcal{T}$  in the standard QFT and the questions related to the status of $\mathcal{C}\mathcal{P}\mathcal{T}$  in BHs physics which were consistently asked by Hawking and ’tHooft seminal works in the last decades. This has caught our attention in building QFT with special emphasis on discrete space-time transformations, which are vital in quantum theory. Acknowledging the fact that time is a parameter in quantum theory, in this paper, we formulated a direct-sum QFT in the context of BH space-time based on our previous study in the context of Minkowski, de Sitter, and inflationary space-times [^9] The framework of DQFT for inflationary quantum fluctuations is found to explain long-standing CMB anomalies [31] 650 times better than the standard theory of inflation. Thus, our framework of DQFT passes the needed observational test and can potentially lead to new unknown effects in BH physics. [15, 31].
Our prescription of quantization together with ’tHooft gravitational backreaction effects has revealed a new structure for understanding quantum fields in BH space-time. Note that all our study is restricted to Schwarzschild BH in this paper. Our BH DQFT implies no entanglement between interior and exterior Hawking particles pairs, at the same time we obtained that the exterior radiation is not independent of the interior due to the new set of non-commutative relations originated from ’tHooft gravitational backreaction effects. This effect is absent from Hawking’s original 1975 paper. The prescription of using direct-sum Fock space and the associated field operators has revealed that quantum fields evolving forward in time would end up as interior states, and those evolving backward in time would lead to the evaporation of the BH. We have discussed in detail the implications of our result for further explorations in treating quantum fields in collapse and evaporation of BH and future research in quantum gravity.
Acknowledgments
KSK was supported by JSPS and KAKENHI Grant-in-Aid for Scientific Research No. JP20F20320 and No. JP21H00069. KSK would like to thank the Royal Society for supporting in the name of Newton International Fellowship. KSK would like to thank Mainz U. and Beira Interior U. for their hospitality, which is where part of the work has been carried out. J. Marto is supported by the grant UIDB/MAT/00212/2020. We want to thank Chris Ripken for the initial collaboration on the project and for giving us a lot of insights into the algebraic QFT. We thank Gerard ’tHooft for truly inspiring us over the years with his impactful insights into the subject. Also, we thank him for the very useful comments and discussions while preparing this manuscript. We thank Yashar Akrami, Norma G. Sanchez, Masahide Yamaguchi, Alexei A. Starobinsky, Luca Buoninfante, Francesco Di Filippo, Yasha Neiman, Paolo Gondolo, Martin Reuter, Paulo V. Moniz, Gia D’vali, Mathew Hull, Enrique Gaztañaga for useful discussions. KSK would like to especially thank Sivasudhan Ratnachalam for his positive encouragement, friendship, and useful discussions on quantum mechanics.
Appendix A DQFT in Rindler spacetime
In this section, we briefly review the formulation of DQFT in Rindler spacetime, which worked out in [21]. This section will give a complementary understanding of DQFT in BH spacetime with the DQFT in Rindler spacetime Sec. 4.2.
Rindler spacetime is the coordinate transformation of Minkowski spacetime. Here, we consider (1+1) dimension Minkowski spacetime represented by
$ds^{2}=-dt^{2}+dz^{2}$ (A.1)
$z^{2}-t^{2}$ $=\frac{1}{a^{2}}e^{2a\xi}\implies\begin{cases}z=\frac{1}{a}e^{a\xi}\cosh{a\eta},\quad t=\frac{1}{a}e^{a\xi}\sinh{a\eta}\quad\left(\textrm{Right Rindler}\right)\\ z=-\frac{1}{a}e^{a\xi}\cosh{a\eta},\quad t=\frac{1}{a}e^{a\xi}\sinh{a\eta}\quad\left(\textrm{Left Rindler}\right)\end{cases}$ (A.2)
$\implies\boxed{ds^{2}=e^{2a\xi}\left(-d\eta^{2}+d\xi^{2}\right)}$
$t^{2}-z^{2}$ $=\frac{1}{a^{2}}e^{2a\eta}\implies\begin{cases}t=\frac{1}{a}e^{a\eta}\cosh{a\xi},\quad z=\frac{1}{a}e^{a\eta}\sinh{a\xi}\quad\left(\textrm{Future Kasner}\right)\\ t=-\frac{1}{a}e^{a\eta}\cosh{a\xi},\quad z=\frac{1}{a}e^{a\eta}\sinh{a\xi}\quad\left(\textrm{Past Kasner}\right)\end{cases}$
$\implies\boxed{ds^{2}=e^{2a\eta}\left(-d\eta^{2}+d\xi^{2}\right)}$
The description of this spacetime is presented pictorially in Fig. 5.
Figure 5: In this picture, we depict the Rindler spacetime with the Left, Right ($z^{2}\gtrsim t^{2}$), and Future, Past ($t^{2}\gtrsim z^{2}$) regions in Minkowski plane $\left(t,\,z\right)$. The curved lines with arrows in the Left and Right regions depict a constant acceleration $ae^{-a\xi}$ in the direction $\eta:\infty\to-\infty$ and $\eta:-\infty\to\infty$ respectively. Future and Past Rindler wedges are the degenerate Kasner Universes where the arrows are changing $z:\mp\infty\to\pm\infty$, which means $\eta:\mp\infty\to\pm\infty$. The fuzzy colored lines indicate Rindler Horizons for Left (Yellow), Right (Green), Future (Cyan), and Past (Pink).
The Rindler coordinates can be rewritten as
$U$ $=-\frac{1}{a}e^{-au}<0,\quad$ $V=\frac{1}{a}e^{-av}>0\quad$ $(\rm Right\,Rindler)$ (A.3)
$U$ $=\frac{1}{a}e^{-au}>0,\quad$ $V=-\frac{1}{a}e^{-av}<0\quad$ $(\rm Left\,Rindler)$
$U$ $=\frac{1}{a}e^{-au}>0,\quad$ $V=\frac{1}{a}e^{-av}>0\quad$ $(\rm Future\,Kasner)$
$U$ $=-\frac{1}{a}e^{-au}<0,\quad$ $V=-\frac{1}{a}e^{-av}<0\quad$ $(\rm Past\,Kasner)$
where
$u$ $=\eta-\xi,\quad v=\eta+\xi$ (A.4)
$U$ $=t-z,\quad V=t+z$
These coordinates (A.3) define the Left, Right, Future, and Past Rindler regions separated by horizons carried through discrete transformations on $\left(U,\,V\right)$. These are analogous to the Kruskal coordinates of SBH in (3.13) and (3.16). Comparing Figs. 4 and 5, we can draw similarities between DQFT in BH and Rindler spacetimes.
According to DQFT in Rindler spacetime [21], a (maximally) entangled pure state (for $z^{2}-t^{2}\gtrsim 0$) is a direct-sum of two pure state components in Left Rindler and Right Rindler regions
$|\psi_{LR}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}|\phi_{R1}\rangle\otimes|\phi_{R2}\rangle\\ |\phi_{L1}\rangle\otimes|\phi_{L2}\rangle\end{pmatrix}$ (A.5)
Similarly, one can also deduce the same for Future and Past regions.